OpenStudy (221emily):

I am struggling to complete this problem. I went to class but I didn't quite pay that much attention how to do this particular kind of problem... can someone help please.

OpenStudy (kc_kennylau):

can you just say the problem lol

OpenStudy (221emily):

Suppose y varies directly with x, and y=25 when x=140. What is the value of x when y=36? a.) 140 b.) 165 c.) 201.6 d.)176 I know that you are supposed to use y/x but I just can't remember

OpenStudy (kc_kennylau):

the ratio is y:x=25:140

hartnn (hartnn):

yes when y varies directly with x, y/x = k = constant

OpenStudy (kc_kennylau):

so when y increases to 36, it's actually (36/25) times 25. therefore x becomes (36/25) times 140.

hartnn (hartnn):

so when y=25 , x= 140 first find the value of k, can you ?

OpenStudy (221emily):

0.178571428 I think

hartnn (hartnn):

the value for k ? you can keep k as 25/140 too so, when y = 36 y/x = k 36/x = 25/140 can you find x from here ?

OpenStudy (221emily):

151 ?

hartnn (hartnn):

no...how you got 151 ? 36/x = 25/140 36 = 25x/140 36*140 = 25 x x= 36*140/25 =...?

OpenStudy (221emily):

201.6

OpenStudy (221emily):

I think I got it thank you :)

hartnn (hartnn):

yes correct :) welcome ^_^